if cos3A+cos3B+cos3C=1 then which type of triangle is it?
A)obtuse
B)acute
c)right angled triangle
D)none of these

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

COS3B+COS3C+COS3A=1
COS3B+COS3C+COS3A-1=0
COS3(B+C)/2.COS3(B-C)/2-SIN^3A/2=0
SOLVING WE GET
COS3C/2.COS3B/2.COS3A/2=0
EITHER3C/2=
C=/3     THEREFORE TRIANGLE IS ACUTE ANGLED

Harditya
9 years ago
The last part is wrong! 3c/2=pi
c=2pi/3
therefore triangle is obtuse